# Natural Language Processing for Dialects of a Language: A Survey

### v2

@@ Abstract.
State-of-the-art natural language processing (NLP) models are trained on massive training corpora, and report a superlative performance on evaluation datasets. This survey delves into an important attribute of these datasets: the dialect of a language. Motivated by the performance degradation of NLP models for dialectic datasets and its implications for the equity of language technologies, we survey past research in NLP for dialects in terms of datasets, and approaches. We describe a wide range of NLP tasks in terms of two categories: natural language understanding (NLU) (for tasks such as dialect classification, sentiment analysis, parsing, and NLU benchmarks) and natural language generation (NLG) (for summarisation, machine translation, and dialogue systems). The survey is also broad in its coverage of languages which include English, Arabic, German among others. We observe that past work in NLP concerning dialects goes deeper than mere dialect classification, and . This includes early approaches that used sentence transduction that lead to the recent approaches that integrate hypernetworks into LoRA. We expect that this survey will be useful to NLP researchers interested in building equitable language technologies by rethinking LLM benchmarks and model architectures.
Blue text macro. Red text macro. Black text macro.
@@ 1. Introduction
Natural language processing (NLP) is an area of artificial intelligence that deals with the processing of text. NLP tasks are broadly viewed as two categories: natural language understanding (NLU) and natural language generation (NLG). The former covers tasks such as sentiment analysis, where the input is a textual sequence, but the output is a label or a set of labels. The latter includes tasks such as summarisation, where both the input and the output are textual sequences. The state-of-the-art in NLP, for both NLU and NLG, is based on Transformer-based models. Large language models (LLMs) that use decoders in the Transformer architecture have significantly increased attention toward NLP from several domains, including but not limited to science, engineering, and technology. LLMs released by commercial organisations report an increasingly higher number of parameters and, resultantly, improved performances on several NLP tasks on benchmark datasets centered around tasks. NLP approaches using LLMs are largely viewed as black-box models trained on massive corpora whose composition is not accurately known. This survey dissects one of many attributes in which variations may exist in the training and test corpora: dialects of a language. Therefore, this survey presents NLP techniques for dialects of a language.
@@
A dialect is defined as the regionally or locally based varieties of a language (Haugen, 1966). In general, a dialect combines the structure and vocabulary of two or more languages, similar to Creole languages [^1] A Creole language, or simply creole, is a stable natural language that develops from the process of different languages simplifying and mixing into a new form (often, a pidgin), and then that form expanding and elaborating into a full-fledged language with native speakers, all within a fairly brief period.222 (Lent et al., 2023) is a recent work that highlights social and scholarly stigmatisation of Creole languages that has resulted in limited advances in NLP for these languages. We also acknowledge the association between perceived social hierarchies and dialects of a language (Kroch, 1986) or are referred to as national variants of a language. A unique example is Australian English which itself derives its phonemes from Southern British English (Cox and Palethorpe, 2007; Cox, 2006) but has also developed its own unique vocabulary (Moore, 1999). In this paper, we use the word ‘dialect’ interchangeably with ‘national/cultural variants’ of a language while acknowledging implications including the perpetuation of social hierarchies. The focus of this survey is the adaptation and evaluation of NLP tasks and approaches on dialects of a language.
@@
In general, our survey is catalysed by the recent efforts in extending LLMs on NLP tasks for dialects of different languages. As researchers continue to look ‘under the hood’ of LLMs, we believe that dialectic differences in training and testing datasets will become increasingly scrutinised, and models will be enhanced to become robust to dialects. As a result, we hope that this survey will help readers and researchers understand past work in NLP techniques for dialects of a language, and contribute to ideas about fair and equitable NLP in the future. Our survey covers a broad range of approaches from classical NLP as well as modern NLP where pre-trained language models (PTLMs) are leveraged.
@@
There have been surveys in the past in overlapping areas.  (Zampieri et al., 2020) is a survey of NLP for dialects spanning languages from around the world. The paper describes in detail the available corpora, and then summarises past approaches to fundamental NLP problems such as POS tagging and parsing, along with applications to NLP. Our survey builds upon theirs in three ways. Firstly, we cover a wider range of downstream tasks, namely summarisation, sentiment analysis and so on. Also, this survey contains recent papers, which highlight increasingly growing attention towards NLP for dialects. Finally, the exposition of our survey follows a deep learning-centric view: we divide past work into NLU and NLG. Another survey by (Blodgett et al., 2020) describes biases of different kinds in an analysis of language technologies, including dialectic bias. We derive from their survey to formulate the motivation and trends in NLP for dialects. Finally, extensive surveys focussing on languages from the Middle East have been reported (Darwish et al., 2021; Shoufan and Alameri, 2015). These are surveys of NLP for standard and dialectic Arabic, primarily focusing on dialect identification and synthesis in the form of machine translation. Our survey unifies the efforts in dialects of languages belonging to multiple language families. The contribution of our survey is:
@@
  • We present past work in terms of NLU and NLG tasks.
@@
  • We highlight trends and future directions, and provide summary tables that will help researchers interested in dialectic NLP research.
@@
  • The survey connects past research with cutting-edge research using state-of-the-art models.
The rest of the paper is organised as follows. We motivate the need for a discussion on dialects in Section 2. We define the scope of the paper and highlight key trends in Section 3. We then cover dialect-specific resources in Section 4. Following that, Section 5 covers several NLU tasks: dialect identification, sentiment analysis, parsing, and NLU benchmarks. Section 6 presents relevant approaches in NLG for machine translation, summarisation and so on. Finally, we conclude the survey and discuss future work in the context of NLP research as well as social/ethical implications in Section 7. The survey contains several summary tables that will be useful for future research in the area.
@@ 2. Motivation
We now motivate the need for this survey via four aspects. We first describe a set of linguistic studies that study dialectic variations to highlight the relevant NLP challenges. We then show why a deeper analysis in dialects is important to better benchmark LLMs, and NLP techniques and how it has ramifications for the responsible and equitable development of technologies. Finally, recent papers focusing on the dialect-robust development of NLP techniques indicate an emerging trend in NLP research. The four aspects highlight that the survey will provide pointers to the future of linguistic and cultural inclusion in NLP.
@@ 2.1. Linguistic Challenges Posed by Dialects
Linguistic researchers have studied the characteristics of non-native users of a language. (Nagata, 2014) study the influence of native language on a person’s usage of English. With the help of native speakers of Asian languages, they show that the probability of native language influence is invariant with English language proficiency. (Lee and Jun, 2016) analyse four dialects of English (British, Indian, Philippines, and American) and show linguistic variations between them. They particularly highlight different preferences for modals among the four dialects. (Haugh and Bousfield, 2012) describe the difference between Australian and English data stating: “The focus of jocular mockery and abuse about the job or profession of the target involved their (attested) lack of ability or skill in his chosen area or profession.”. This holds true in several other cultures (Haugh and Chang, 2015) including Chinese (Qiu et al., 2021), Indian (Kachru, 1965) and Persian (Izadi, 2015).
Dialectic differences primarily occur in terms of syntax and vocabulary. Some examples of dialectic differences in English are: ‘I might could help you with that’ in Australian English (Coats, 2022), ‘Inside tent can not see leh !’ in Singaporean English (Wang et al., 2017). or the arbitrary placement of adverbs in native speakers of Asian languages as in ‘Already, I have done it.’ (Nagata, 2014). Also, consider the case of the Samvedi [^3] Samvedi is also called Kadodi and is primarily spoken by the Samvedi Brahmins and Catholics predominantly residing in Vasai City, Maharashtra, India dialect of Marathi, one of 42444https://en.wikipedia.org/wiki/Marathi_language#Dialects, where the Samvedi version of the modern standard Marathi sentence “(Transliteration) Mazha naav Pushpak ahe”(Literal: My name Pushpak is) (Translation: My name is Pushpak) is “ (Transliteration) Maa naav Pushpak hai”(Literal: My name Pushpak is) (Translation: My name is Pushpak). Samvedi does not exhibit word order differences compared to standard Marathi, but it involves heavy pronunciation relaxation (ahe -¿ ha, and mazha-¿maa, in the example above) and the usage of older words.
Another challenge in handling dialects is that two dialects of the same language can be completely mutually unintelligible, and is often correlated with the distance between regions in which that dialect is primarily spoken. A classic example of this is in the case of the Aomori [^5] https://en.wikipedia.org/wiki/Tsugaru_dialect and Okinawan [^6] https://en.wikipedia.org/wiki/Okinawan_Japanese dialects of Japanese which has a total of 47777https://en.wikipedia.org/wiki/Japanese_dialects known dialects. Therefore, it is not enough to collect data for one dialect and assume that it will help in NLP for another dialect which indicates that special attention will need to be paid to each dialect to ensure that it will be well-represented.
Dialects assume further importance when people from different cultural backgrounds interact with each other. (Meyer, 2014) discusses interactions between people of different cultural backgrounds. The book compares interactions of Australians with people from other cultures in terms of (a) building trust with colleagues, (b) leading teams of a culturally dissimilar background etc. An example in the book states that an Australian may invest in shorter small talk than a Mexican with a colleague. (Wang et al., 2022) show that monophthongal vowels spoken by Australian English speakers may be difficult to be understood by Mandarin English listeners.
@@
Further challenges are posed by differences in pragmatics, with influences derived from macro-social factors such as region, social class, ethnicity, gender, age (Haugh and Schneider, 2012). For example, (Schneider, 2012) observes differences in small talk across inner circle varieties of English which is also linked to age and gender. (Merrison et al., 2012) showed that in student requests to university staff, there were differences in the way obligation was expressed, and that these differences were linked to different ways of claiming social standing.
Finally, it is important to acknowledge that a large proportion of users of English are L2 speakers, are using English not as a second language in everyday life but have acquired English in school, and use it as a tool for communication when required, e.g., in business settings or on holiday. Research in inter-language pragmatics (see summary (Kecskés, 2022) has also shown that there are systematic differences in how L2 learners differ use pragmatic strategies, and how different factors such as length of exposure, study abroad, proficiency) influence these strategies.
Noting the differences in the pragmatic strategies of different dialect speakers provide an important social perspective on dialectal variation. However, these are currently not sufficiently accounted for in NLP.
@@ 2.2. Rethinking LLM benchmarks
Taking the example of English, there are more English language speakers in countries such as India than the United States, Australia and England (Dunn, 2019). In addition, an even larger number of speakers have acquired English in a classroom context (e.g., in countries such as China, Germany or Russia) and use it mainly as a contact language for specific transactional purposes, e.g., business or education. This latter perspective has been described through the notion of English as a lingua franca as “the common language of choice […] among speakers who come from different lingua-cultural backgrounds” (Jenkins, 2009). As (Seidlhofer, 2005) points out, “English is being shaped at least as much by its non-native speakers as by its native speakers”.
Yet, the corpora used to train language models and more importantly, the datasets used to evaluate them do not necessarily reflect this gap. (Inoue et al., 2021) examine the performance of BERT-based models for varieties/dialects of Arabic, and show that dialect proximity of pre-training and fine-tuning data bears impact on the performance of the downstream task. In the case of GPT-4, the evaluation dataset consists of questions from the MMLU benchmark written in Standard American English [^8] GPT-4 also claims multilingual performance in various languages using translations of English questions into target languages. Using translations of sentences to claim multilingual ability is inaccurate (Joshi et al., 2023). Standard benchmarks used to claim performance of a language model for English primarily contain Standard American English. It has been found that the performance does not extend to NLU tasks for dialects of English (Ziems et al., 2022). This holds for most foundation models that are trained on large amounts of data. The distribution of languages in the training corpora is either not known or difficult to determine. There are variants of English, i.e., dialects, that these models brush under the carpet, despite the existence of corpora of regional varieties of English [^9] https://www.ice-corpora.uzh.ch/en.html; Accessed on 9th January, 2024. and of corpora of English as a lingua franca [^10] https://www.kielipankki.fi/corpora/elfa/; Accessed on 9th January, 2024.. Some examples showing the impact of dialects on the performance of NLP tasks are presented in Table 1. We note that these papers are from the past few years, which have otherwise witnessed a great development in the reported performance of NLP models.
@@ 2.3. Fair and equitable technologies
In a multicultural society, people use different dialects. NLP systems that are deployed to serve such multicultural communities must be mindful of the variations between different dialects. Evaluation and mitigation of disparity become an overgrowing need in times when language models claim excellent language performance using datasets from a specific dialect alone. This holds particular importance in terms of the fairness of NLP approaches.
The following papers show the implications of dialects along sociological factors:
@@
    Performance of NLP models and per-capita GDP: A recent work by (Kantharuban et al., 2023) show the dialectic gap in performance of LLM-based solutions for machine translation and automatic speech recognition for several dialects including those of Arabic, Mandarin, Finnish, German, Bengali, Tagalog and Portuguese. This shows that the problem is not focussed on a specific region but on languages of the world. The paper also analyses confounding social factors and the associated impact on the size of digitized corpora. They show a positive correlation between gross domestic product per capita and the efficacy of dialectic machine translation.
@@
    Healthcare monitoring(Jurgens et al., 2017) show that there exists a disparity between popular dialect speakers and others in the case of healthcare monitoring [^11] They also propose a method to mitigate the disparity..
@@
    Racial biases in hate speech detection(Okpala et al., 2022) show that hate speech classifiers may lean towards predicting a text as true if it uses African-American English.
@@
    Prejudice in prediction of employability and criminality: (Hofmann et al., 2024) show that dialects may introduce bias in the output of language models. As a result, a person’s output with respect to their employability or criminality may be affected based on the dialects they use.
@@
An idea closely related to the survey is the ‘Bender rule’ in NLP research. The Bender rule states that the language of datasets used for evaluation must be stated explicitly without assuming English to be the implicit default (Ducel et al., 2022). We similarly believe that languages are not monoliths and dialectic differences must be clearly stated.
@@
Parsing  (Scannell, 2020) Lower performance of parsers on Mancks Gaellic
as compared to Irish/Scottish Gaellic.
Table 1. Examples of adverse impact on NLP task performance due to dialectic variations.

### v4

@@ Abstract.
State-of-the-art natural language processing (NLP) models are trained on massive training corpora, and report a superlative performance on evaluation datasets. This survey delves into an important attribute of these datasets: the dialect of a language. Motivated by the performance degradation of NLP models for dialectal datasets and its implications for the equity of language technologies, we survey past research in NLP for dialects in terms of datasets, and approaches. We describe a wide range of NLP tasks in terms of two categories: natural language understanding (NLU) (for tasks such as dialect classification, sentiment analysis, parsing, and NLU benchmarks) and natural language generation (NLG) (for summarisation, machine translation, and dialogue systems). The survey is also broad in its coverage of languages which include English, Arabic, German, among others. We observe that past work in NLP concerning dialects goes deeper than mere dialect classification, and extends to several NLU and NLG tasks. For these tasks, we describe classical machine learning using statistical models, along with the recent deep learning-based approaches based on pre-trained language models. We expect that this survey will be useful to NLP researchers interested in building equitable language technologies by rethinking LLM benchmarks and model architectures.
@@ 1. Introduction
Natural language processing (NLP) is an area of artificial intelligence that deals with processing of human language in its textual form. NLP tasks are broadly viewed as two categories: natural language understanding (NLU) and natural language generation (NLG). The former broadly covers language understanding tasks such as dialect identification or sentiment classification, as well as tasks such as morphosyntactic analysis. The latter includes tasks where both the input and the output are textual sequences (for example, summarisation). The state-of-the-art NLP, for both NLU and NLG, is based on Transformer-based models (Naveed et al., 2023; Zhao et al., 2023). Large language models (LLMs) that use decoders in the Transformer architecture have significantly increased attention toward NLP leading to LLM-based applications in several domains such as medicine, business or law. LLMs released by commercial organisations report an increasingly higher number of parameters and, as a result, improved performances on several NLP tasks. NLP approaches using LLMs are largely viewed as black-box models trained on massive corpora whose composition is not accurately known. This survey dissects one of many attributes in which variations may exist in the training and test corpora: dialects of a language.
@@
Traditionally, a dialect is defined as the regionally or locally based variety of a language (Haugen, 1966). Wikipedia defines a dialect as “a variety of a language that is a characteristic of a particular group of the language’s speakers.” Zampieri and Nakov (2021) state that dialects are language varieties characterised by systematic patterns of variation. The current notion of dialect has extended to language varieties arising due to factors such as political reasons, country of origin, migration histories, historical factors, register shifts and so on. In fact, there is an association between perceived social hierarchies and dialects of a language, leading to a term ‘sociolect’ (Kroch, 1986). For the sake of brevity, we use ‘dialects’ as an umbrella term to refer todialects/national varieties/cultural variants/sociolects’ of a language while acknowledging that the distinction between dialects and language is nuanced (Sandel, 2015). An example of a dialect is the national variety, Australian English, whose phonemes are predominantly derived from Southern British English and other Englishes (Cox and Palethorpe, 2007; Cox, 2006), but has also developed its own unique vocabulary (Moore, 1999). Overlapping with dialects are Creole languages that develop from the process of different languages simplifying and mixing into a new form (often, a pidgin), and then that form expanding and elaborating into a full-fledged language with native speakers, all within a fairly brief period. Lent et al. (2024) highlight the social and scholarly stigmatisation of Creole languages that has resulted in limited advances in NLP for these languages.
@@
In general, our survey is catalysed by the recent efforts in extending LLMs on NLP tasks for dialects of different languages. As researchers continue to look ‘under the hood’ of LLMs, dialectal differences in training and testing datasets are being increasingly scrutinised, and adaptation techniques to improve their performance on different dialects are being devised. As a result, we hope that this survey will help readers and researchers understand past work in NLP techniques for dialects of a language, and contribute to ideas about fair and equitable NLP in the future.
@@
There have been related surveys in the pastZampieri et al. (2020) describes the available corpora, and past approaches to fundamental NLP problems such as POS tagging and parsing, along with applications to NLP. Our survey builds upon theirs in three ways. Firstly, we cover a wider range of downstream tasks such as summarisation and sentiment analysis. Also, this survey contains recent papers, which highlight increasingly growing attention towards NLP for dialects. Finally, the exposition of our survey adopts a deep learning-centric view, by covering deep learning-based approaches, in particular, the recent LLM-based approaches. Another survey by Blodgett et al. (2020) describes biases of different kinds in an analysis of language technologies, including dialectal bias. We derive from their survey to formulate the motivation and trends in NLP for dialects. Similarly, Jauhiainen et al. (2019) present a survey of automatic language identification, which does not differentiate between dialect or language identification, and mention that dialect identification may be a more challenging task. Finally, extensive surveys focusing on languages from the Middle East have been reported (Darwish et al., 2021; Shoufan and Alameri, 2015). These are surveys of NLP for standard and dialectal Arabic, primarily focusing on dialect identification and synthesis in the form of machine translation. Our survey unifies the efforts in dialects of languages belonging to multiple language families. The contribution of our survey is:
@@
  • We present past work in terms of NLU and NLG tasks, and include both pre-deep learning and deep learning techniques.
@@
  • We highlight trends and future directions, and provide summary tables that will help researchers interested in dialectal NLP research.
@@
The rest of the paper is organised as follows. We motivate the need for a discussion on dialects in Section 2. We define the scope of the paper and highlight key trends in Section 3. We then cover dialect-specific resources in Section 4. Following that, Section 5 covers several NLU tasks: dialect identification, sentiment analysis, parsing, and NLU benchmarks. Section 6 presents relevant approaches in NLG for machine translation, summarisation and so on. Finally, we conclude the survey and discuss future work in the context of NLP research as well as social/ethical implications in Section 7. The survey contains several summary tables that will be useful for future research.
@@ 2.1. Linguistic Challenges Posed by Dialects
Figure 1. An example sentence highlighting the differences between Marathi and its Samvedi dialect.
Dialectal differences primarily occur in terms of orthography, syntax and vocabulary. Some examples of dialectal differences in English are: ‘I might could help you with that’ observed in Southern US, Australian and New Zealand English (Morin and Coats, 2023; Coats, 2022) as well as British and Irish English (Coats, 2023), ‘Inside tent can not see leh !’ in Singaporean English (Wang et al., 2017) or the uncommon placement of adverbs in native speakers of Asian languages as in ‘Already, I have done it.’ (Nagata, 2014). Also, consider the case of the Samvedi dialect of Marathi, one of $42$, where we give an example of the Samvedi and Marathi sentences in Table 1. Samvedi does not exhibit word order differences compared to standard Marathi, but it involves heavy pronunciation relaxation (ahe -¿ hay, and maza -¿ maa) and the usage of older words. Another challenge in handling dialects is that two dialects of the same language can be mutually unintelligible. A classic example of this is the case of the Aomori and Okinawan dialects of Japanese, which has a total of 47 known dialects. Therefore, it is not enough to collect data for one dialect and assume that it will help in NLP for another dialect, which indicates that special attention will need to be paid to each dialect to ensure that it will be well-represented. Dialects assume further importance when people from different cultural backgrounds interact with one another.  Wang et al. (2022) show that monophthongal vowels spoken by Australian English speakers may be difficult to be understood by Mandarin English listeners.
@@
Dialects are also associated with pragmatics, with influences derived from macro-social factors such as region, social class, ethnicity, gender, age (Haugh and Schneider, 2012). For example, Schneider (2012) observes differences in small talk across inner circle varieties of English, i.e., varieties from countries where it is the primary language (Kachru, 1992). They also observed differences between speakers of different ages and genders. This suggests that the notion of ‘dialect’ can be linked to factors beyond geographical distribution. Merrison et al. (2012) showed that, in student requests to university staff, there were differences in the way obligation was expressed, and that these differences were linked to different ways of claiming social standing.  Meyer (2014) compares interactions of Australians with people from other cultures in terms of (a) building trust with colleagues, (b) leading teams of a culturally dissimilar background etc. An example in the book states that an Australian may invest in shorter small talk than a Mexican with a colleague. Noting the differences in the pragmatic strategies of different dialect speakers provide an important social perspective on dialectal variation. However, these are currently not sufficiently accounted for in NLP.
@@ 2.2. Rethinking LLM benchmarks
There are more English language speakers in countries such as India than the United States, Australia and England (Dunn, 2019). In addition, an even larger number of speakers have acquired English in a classroom context (e.g., in countries such as China, Germany or Russia) and use it mainly as a contact language for specific transactional purposes, e.g., business or education. This latter perspective has been described through the notion of English as a lingua franca as “the common language of choice […] among speakers who come from different lingua-cultural backgrounds” (Jenkins, 2009). Despite that, the corpora used to train language models and more importantly, the datasets used to evaluate them do not necessarily reflect dialectal variations within a language. Inoue et al. (2021) examine the performance of BERT-based models for varieties/dialects of Arabic, and show that dialect proximity of pre-training and fine-tuning data bears impact on the performance of the downstream task. In the case of GPT-4, the evaluation dataset consists of questions from the MMLU benchmark written in Standard American English. Standard benchmarks used to claim performance of a language model for English primarily contain Standard American English. It has been found that the performance does not extend to NLU tasks for dialects of English (Ziems et al., 2022). Further, a recent work by Fleisig et al. (2024) analyses the output of ChatGPT for varieties of English, and shows that the generated output may be of poorer quality and be prone to stereotyping for non-standard dialects of English. These findings holds for most foundation models that are trained on large amounts of data. The distribution of languages in the training corpora is either not known or difficult to determine.
@@ 2.3. Fair and equitable technologies
NLP systems that are deployed to serve multicultural communities must be mindful of the variations between different dialects. Evaluation and mitigation of disparity between dialects become an overgrowing need in times when language models claim excellent language performance using datasets from a specific dialect alone. Some examples showing the impact of dialects on the performance of NLP tasks are presented in Table 1. We note that these papers are from the past few years, which have otherwise witnessed a great development in the reported performance of NLP models.
Some implications of dialects in terms of sociological factors are:
@@
    Performance of NLP models and per-capita GDP: A recent work by Kantharuban et al. (2023) show the dialectal gap in performance of LLM-based solutions for machine translation and automatic speech recognition for several dialects, similar to  Ahia et al. (2023) who show the same for topologically diverse languages. They show a positive correlation between gross domestic product per capita and the efficacy of dialectal machine translation.
@@
    Healthcare monitoringJurgens et al. (2017) show that there exists a disparity between popular dialect speakers and others in the case of healthcare monitoring [^1] They also propose a method to mitigate the disparity..
@@
    Racial biases in hate speech detectionOkpala et al. (2022) show that hate speech classifiers may lean towards predicting a text as true if it uses African-American English.
@@
    Prejudice in the prediction of employability and criminality: Hofmann et al. (2024) show that dialects may introduce bias in the output of language models. As a result, a person’s output with respect to their employability or criminality may be affected based on the dialects they use.
@@
An idea closely related to the survey is the ‘Bender rule’ in NLP research. The Bender rule states that the language of datasets used for evaluation must be stated explicitly without assuming English to be the implicit default (Ducel et al., 2022). We similarly believe that languages are not monoliths and dialectal differences must be clearly stated. Similarly,  Hovy and Yang (2021) show that incorporating dialectal aspects is closely related to social factors of language. As a result, incorporating an understanding of dialects of a dataset is pivoting for fairer NLP tools.
@@
Parsing  (Scannell, 2020) Lower performance of parsers on Manx Gaelic
as compared to Irish/Scottish Gaelic.
Table 1. Examples of adverse impact on NLP task performance due to dialectal variations.